# Estimation of Tsallis entropy and its applications to goodness-of-fit tests

### v1

@@ Abstract
In this paper, we consider the problem of estimating Tsallis entropy from a given data set. We propose four different estimators for Tsallis entropy measure based on higher-order sample spacings, and then discuss estimation of Tsallis divergence measure. We compare the performance of the proposed estimators by means of bias and mean squared error and also examine their robustness to outliers. Next, we propose a spacings-based estimator for Tsallis entropy under progressive type-II censoring and study its performance using Monte Carlo simulations. Another estimator for Tsallis entropy is proposed using quantile function and its consistency and asymptotic normality are studied, and its performance is evaluated through Monte Carlo simulations. Goodness-of-fit tests for normal and exponential distributions as applications are developed using Tsallis divergence measure. The performance of the proposed tests are then compared with some known tests using simulations and it is shown that the proposed tests perform very well. Also, an exponentiality test under progressive type-II censoring is proposed, its performance is compared with existing entropy-based tests using simulation. It is observed that the proposed test performs well. Finally, some real data sets are analysed for illustrative purposes.
I Introduction
@@
Estimation of entropy from a data set is an important problem and it has been addressed by many authors over the years. Since the estimation of Shannon entropy by Vasicek (1976), various estimators have been introduced in the literature; see, for example, Van Es (1992), Ebrahimi et al. (1994), Correa (1995), Noughabi (2010), Al-Omari (2014), and Noughabi and Arghami (2022). Although estimation of entropy has gained a lot of attention, estimation of Rényi and Tsallis entropies have not been studied to the same extent. Hegde et al. (2005) proposed order statistics based $m$-spacings estimator for Rényi entropy and recently, Al-Labadi et al. (2025) introduced an estimator for Rényi entropy motivated by the works of Vasicek (1976) and Ebrahimi et al. (1994). Kernel-based estimation of Tsallis entropy has been studied recently for $\rho$-mixing dependent data and length-biased data; see Maya et al. (2024) and Zamini et al. (2024). To the best of our knowledge, order statistics based non-parametric estimators for Tsallis entropy have not been considered in the literature yet. In this paper, we propose some non-parametric estimators for Tsallis entropy using order statistics and compare their performance in terms of bias and mean squarred error. Also, we propose an estimator for Tsallis divergence measure. An estimator for Tsallis entropy under progressive type-II censored data is also developed and its performance is examined through Monte Carlo simulations. Another estimator for Tsallis entropy based on the derivative of the quantile function (qf) of the underlying rv is introduced and its asymptotic properties are studied. Using the estimator for Tsallis divergence, we also develop goodness-of-fit tests for normal and exponential distributions.
The rest of the paper is organized as follows. We introduce order statistics based non-parametric estimators for Tsallis entropy in Section II. Estimator for Tsallis divergence is studied next in Section III. The performance of these estimators and their robustness to outliers are compared through Monte Carlo simulations in Section IV. An estimator for Tsallis entropy under progressive type-II censoring is discussed in Section V. The quantile-based estimator is studied in Section VI and its performance is investigated in Section VII. Applications to some real data sets are discussed in Section VIII wherein goodness-of-fit tests for normal and exponential distributions are explored for complete and progressively type-II censored samples. Finally, some brief concluding remarks are made in Section IX.
II Estimators for Tsallis entropy based on $m$-spacings
@@
III Estimation of Tsallis divergence
@@
TABLE I: Bias and MSE of $T_{\alpha}V$, $T_{\alpha}H$, $T_{\alpha}E$ and $T_{\alpha}W$ for N(0,1) distribution.
@@
TABLE II: Bias and MSE of $T_{\alpha}V$, $T_{\alpha}H$, $T_{\alpha}E$ and $T_{\alpha}W$ for Exp(1) distribution.
@@ IV Comparison of estimators
The true value of $T_{\alpha}(X)$ for Normal and Exponential distributions are $\frac{1}{\alpha-1}\left(1-\frac{1}{(2\pi\sigma^{2})^{\frac{\alpha-1}{2}}\alpha^{\frac{1}{2}}}\right)$ and $\frac{1}{\alpha-1}\left(1-\frac{\lambda^{\alpha-1}}{\alpha}\right)$, respectively. We calculate the bias and the MSE for the estimators for different choices of $n$, $m$ and $\alpha$. We take $n$ = 20, 50 and 100 and $\alpha$ = 0.5, 1, 1.5, 2, 2.5 and 3. For the choice of $m$, we choose it as $m=[\sqrt{n}+0.5]$, which has been used by Wieczorkowski and Grzegorzewski (1999). Another choice of $m$ is $m=\left[\frac{m}{3}\right]$. So, the choices for $m$ for $n$ = 20, 50 and 100 are (4,7,10) and (6,16,33), respectively. We have presented the obtained results in Tables I and II. From these tables, we observe that the performance of the estimators depend on the choice of $m$. As the sample size increases, bias and MSE of the estimators decrease, as expected. From Table I, we observe that for N(0,1) distribution, $T_{\alpha}W$ outperforms the other estimators when $n$ = 20 for most of the cases. However, $T_{\alpha}V$ performs better than all the estimators in most of the cases for $n$ = 50 and 100. When $\alpha\to 1$, i.e., for the KL divergence estimation, $T_{\alpha}W$ performs the best for small sample sizes and $T_{\alpha}E$ performs the best for large sample sizes. Also, in most of the cases, $T_{\alpha}W$ performs better than $T_{\alpha}E$ and that the estimators perform better for $m=\left[\frac{m}{3}\right]$ than for $m=[\sqrt{n}+0.5]$. In general, for N(0,1) distribution, Vasicek-type estimators are better suited with $m=\left[\frac{m}{3}\right]$. From Table II, it is evident that for Exp(1) distribution, $T_{\alpha}H$ has better performance in most of the cases. However, when the sample size increases, performance of $T_{\alpha}V$ and $T_{\alpha}H$ become similar. Note that $T_{\alpha}E$ has better performance for Exp(1) distribution than N(0,1) distribution. Here, both choices of $m$ are seen to be satisfactory. For $\alpha>1$, $m=\left[\frac{m}{3}\right]$ has overall better performance and the reverse is true for $\alpha$ = 0.5 and 1. From these observations, we can conclude that for practical purposes, Vasicek-type estimator would be quite suitable for both normal and exponential distributions and for both choices of $m$.
@@ IV-A Robustness of the estimators in the presence of outliers
V Estimation of Tsallis entropy based on progressive type-II censored data using $m$-spacings
@@
V-A Empirical evaluation
We perform Monte Carlo simulations to assess the performance of the proposed estimator by examining its mean and variance. We consider different progressive censoring schemes for $n$ = 20 and $r$ = 10 and 15. We choose window size $m=[\sqrt{r}+0.5]$ and $\alpha$ = 1.10, 1.5 and 2. We mostly consider one-step progressive censoring schemes where all the surviving units are removed at one time during the experiment. We simulate 5000 progressively type-II censored samples from each of Exp(1) and N(0,1) distributions and calculate the average estimate (AE) and the variance of $\hat{T_{\alpha}}$. The obtained results are presented in Tables III and IV, respectively. In these tables, the notation aba*b refers to ‘$a$ is repeated $b$ times’. For instance, (05,10,04)(0*5,10,0*4) refers to $(0,0,0,0,0,10,0,0,0,0)$.
TABLE III: Average Estimate and Variance of $\hat{T_{\alpha}}$ for Exp(1) distribution.
@@ V-A Empirical evaluation
TABLE IV: Average Estimate and Variance of $\hat{T_{\alpha}}$ for N(0,1) distribution.
@@
VI Estimation of Tsallis entropy through quantile approach
@@ 3. $K(\cdot)$ is symmetric about zero;
Proceeding along the same lines as in Theorems VI.1 and VI.2, we can establish the consistency and asymptotic normality of $\tilde{T}_{\alpha}(X)$ as well.
VII Empirical study
@@
and is denoted by Govindarajulu ($\mu,\sigma,\gamma$). We generate 5000 random samples from N(0,1) and Exp(1) distributions with sample size $n$ = 20, 50 and 100. We consider $\alpha$ = 0.5, 1.5, 2, 2.5 and 3 and then calculate the bias and the MSE of $T^{q}_{\alpha}(X)$. We use the Gaussian kernel and select bandwidth according to the normal reference rule $h_{n}=1.06\sigma n^{-\frac{1}{5}}$ (cf. Silverman, 2018). The obtained results are provided in Table V. From this table, we observe that as the sample size increases, the bias and the MSE of $T^{q}_{\alpha}(X)$ for both distributions decrease. The estimator performs better when $\alpha$ increases. Also, $T^{q}_{\alpha}(X)$ performs better for N(0,1) than Exp(1) distribution. This may be due to the fact that we have chosen Gaussian Kernel and that the choice of $h_{n}$ is appropriate when data come from normal distribution.
TABLE V: Bias and MSEs of $T^{q}_{\alpha}(X)$ for N(0,1) and Exp(1) distributions.
@@
There are some distributions that do not have tractable cdf, but have simple and closed form of qf, such as power-Pareto, skew lambda and Govindarajulu distributions. In these situations, it is advantageous to use $T^{q}_{\alpha}(X)$ to estimate Tsallis entropy. We simulate 5000 random samples from Govindarajulu distribution with parameters (0,0.75,0.25) and (0,0.25,0.75), for $n$ = 50, 100 and 200, and then calculate the bias and the MSE of $T^{q}_{\alpha}(X)$. We chose $\alpha$ = 0.15 and 0.75. Note that, for these models, Tsallis entropy does not exist for $\alpha>1$. We provide the obtained results in Table VI. The true value of Tsallis entropy are also included in the table. From this table, we observe that the estimator performs well for these models. Both Bias and MSE decrease as sample size increases and the performance of the estimator seems to be better when $\alpha$ is small.
TABLE VI: Bias and MSEs of $T^{q}_{\alpha}(X)$ for Govindarajulu distribution.
@@
VIII Applications to Goodness-of-fit testing
@@
TABLE VII: Power of the tests for N(0,1) distribution.
@@ VIII-A Test for Normality
TABLE VIII: Power of the tests for N(0,1) distribution.
@@ VIII-B Test for Exponentiality
The null hypothesis is rejected when the calculated value of $T>C_{2}$, where the critical point $C_{2}$ is obtained from the null distribution of $T$. For reference distribution, we consider Weibull (WE), gamma (GA) and log-normal (LN) distributions having monotone decreasing, increasing and non-monotone hazard functions. For monotone decreasing hazard alternatives, we take WE(0.5), GA(0.4) and LN(2). For monotone increasing hazard alternatives, we consider WE(2), GA(2) and GA(3), while for non-monotone hazard alternatives, we consider LN(0.6) and LN(1.2). The power of the tests so obtained are provided in Tables IX and XI. From these tables, we find that proposed test performs better than the other two tests for monotone decreasing hazard alternative distributions.
TABLE IX: Power of the tests for Exp(1) distribution.
@@
VIII-C Goodness-of-fit Tests for Progressive Type-II Samples
@@
TABLE X: Power of exponentiality tests under PC-II samples at 10% significance level.
@@
VIII-C1 Exponentiality under Progressive Type-II Samples
For exponentiality testing, the test statistic is given by
where $\hat{\theta}=\frac{r}{\sum_{i=1}^{r}(1+R_{i})X_{i:r:n}}$ is the MLE for $\theta$ under PC-II censoring. We will reject exponentiality hypothesis if $\hat{T_{\alpha}}(F_{r,n},F)$ is large. We simulate 5000 random samples from Exp(1) distributions under various PC-II censoring schemes and obtain critical points at 10% significance level. Entropy based exponentiality tests under PC-II censoring perform better when the alternative distribution has increasing hazard function, see Balakrishnan et al. (2007). So, we consider WE(2) and GA(2) distributions for power calculation and compare the performance of the proposed test with the test proposed by Balakrishnan et al. (2007). Let us denote it by $T_{PC}$. We choose $\alpha$ = 2 and $m=[\sqrt{r}+0.5]$ for power computation and provide the results in Table X. From the table we have observed that neither test outperforms the other for all the cases. Performance of both tests increases when sample size increases. Both tests perform better for WE(2) than for GA(2) alternative. The proposed test does attain the nominal level of significance. We mostly considered one step censoring schemes for illustrations. For other censoring schemes, the proposed tests outperforms $T_{PC}$ for both the alternative distributions.
VIII-D Data Analyses

### v2

@@ Abstract
In this paper, we consider the problem of estimating Tsallis entropy from a given data set. We propose four different estimators for Tsallis entropy measure based on higher-order sample spacings, and then discuss estimation of Tsallis divergence measure. We compare the performance of the proposed estimators by means of bias and mean squared error and also examine their robustness to outliers. Next, we propose a spacings-based estimator for Tsallis entropy under progressive type-II censoring and study its performance using Monte Carlo simulations. Another estimator for Tsallis entropy is proposed using quantile function and its consistency and asymptotic normality are studied, and its performance is evaluated through Monte Carlo simulations. Goodness-of-fit tests for normal and exponential distributions as applications are developed using Tsallis divergence measure. The performance of the proposed tests are then compared with some known tests using simulations and it is shown that the proposed tests perform very well. Also, an exponentiality test under progressive type-II censoring is proposed, its performance is compared with an existing entropy-based test using simulation. It is observed that the proposed test performs well. Finally, some real data sets are analysed for illustrative purposes.
Keywords and Phrases: Empirical influence function, goodness-of-fit tests, Kullback-Leibler divergence, order statistics, power analysis, progressive censoring, quantile density function
Mathematics Subject Classification–62G05, 62N02, 94A17
1 Introduction
@@
Estimation of entropy from a data set is an important problem and it has been addressed by many authors over the years. Since the estimation of Shannon entropy by Vasicek (1976), various estimators have been introduced in the literature; see, for example, Van Es (1992), Ebrahimi et al. (1994), Correa (1995), Noughabi (2010), Al-Omari (2014), and Noughabi and Arghami (2022). Although estimation of entropy has gained a lot of attention, estimation of Rényi and Tsallis entropies have not been studied to the same extent. Hegde et al. (2005) proposed order statistics based $m$-spacings estimator for Rényi entropy and recently, Al-Labadi et al. (2025) introduced an estimator for Rényi entropy motivated by the works of Vasicek (1976) and Ebrahimi et al. (1994). Kernel-based estimation of Tsallis entropy has been studied recently for $\rho$-mixing dependent data and length-biased data; see Maya et al. (2024) and Zamini et al. (2024). To the best of our knowledge, order statistics based non-parametric estimators for Tsallis entropy have not been considered in the literature yet. In this paper, we propose some non-parametric estimators for Tsallis entropy using order statistics and compare their performance in terms of bias and mean squarred error. Also, we propose an estimator for Tsallis divergence measure. An estimator for Tsallis entropy under progressive type-II censored data is also developed and its performance is examined through Monte Carlo simulations. Another estimator for Tsallis entropy based on the derivative of the quantile function (qf) of the underlying rv is introduced and its asymptotic properties are studied. Using the estimator for Tsallis divergence, we also develop goodness-of-fit tests for normal and exponential distributions. Goodness-of-fit test for exponential distribution under progressive type-II censored sample is also discussed.
The rest of the paper is organized as follows. We introduce order statistics based non-parametric estimators for Tsallis entropy in Section 2. Estimator for Tsallis divergence is studied next in Section 3. The performance of these estimators and their robustness to outliers are compared through Monte Carlo simulations in Section 4. An estimator for Tsallis entropy under progressive type-II censoring is discussed in Section 5. The quantile-based estimator is studied in Section 6 and its performance is investigated in Section 7. Applications to some real data sets are discussed in Section 8 wherein goodness-of-fit tests for normal and exponential distributions are explored for complete and progressively type-II censored samples. Finally, some brief concluding remarks are made in Section 9.
2 Estimators for Tsallis entropy based on $m$-spacings
@@
3 Estimation of Tsallis divergence
@@
Table 1: Bias and MSE of $T_{\alpha}V$, $T_{\alpha}H$, $T_{\alpha}E$ and $T_{\alpha}W$ for N(0,1) distribution.
@@
Table 2: Bias and MSE of $T_{\alpha}V$, $T_{\alpha}H$, $T_{\alpha}E$ and $T_{\alpha}W$ for Exp(1) distribution.
@@ 4 Comparison of estimators
The true value of $T_{\alpha}(X)$ for Normal and Exponential distributions are $\frac{1}{\alpha-1}\left(1-\frac{1}{(2\pi\sigma^{2})^{\frac{\alpha-1}{2}}\alpha^{\frac{1}{2}}}\right)$ and $\frac{1}{\alpha-1}\left(1-\frac{\lambda^{\alpha-1}}{\alpha}\right)$, respectively. We calculate the bias and the MSE for the estimators for different choices of $n$, $m$ and $\alpha$. We take $n$ = 20, 50 and 100 and $\alpha$ = 0.5, 1, 1.5, 2, 2.5 and 3. For the choice of $m$, we choose it as $m=[\sqrt{n}+0.5]$, which has been used by Wieczorkowski and Grzegorzewski (1999). Another choice of $m$ is $m=\left[\frac{m}{3}\right]$. So, the choices for $m$ for $n$ = 20, 50 and 100 are (4,7,10) and (6,16,33), respectively. We have presented the obtained results in Tables 1 and 2. From these tables, we observe that the performance of the estimators depend on the choice of $m$. As the sample size increases, bias and MSE of the estimators decrease, as expected. From Table 1, we observe that for N(0,1) distribution, $T_{\alpha}W$ outperforms the other estimators when $n$ = 20 for most of the cases. However, $T_{\alpha}V$ performs better than all the estimators in most of the cases for $n$ = 50 and 100. When $\alpha\to 1$, i.e., for the KL divergence estimation, $T_{\alpha}W$ performs the best for small sample sizes and $T_{\alpha}E$ performs the best for large sample sizes. Also, in most of the cases, $T_{\alpha}W$ performs better than $T_{\alpha}E$ and that the estimators perform better for $m=\left[\frac{m}{3}\right]$ than for $m=[\sqrt{n}+0.5]$. In general, for N(0,1) distribution, Vasicek-type estimators are better suited with $m=\left[\frac{m}{3}\right]$. From Table 2, it is evident that for Exp(1) distribution, $T_{\alpha}H$ has better performance in most of the cases. However, when the sample size increases, performance of $T_{\alpha}V$ and $T_{\alpha}H$ become similar. Note that $T_{\alpha}E$ has better performance for Exp(1) distribution than N(0,1) distribution. Here, both choices of $m$ are seen to be satisfactory. For $\alpha>1$, $m=\left[\frac{m}{3}\right]$ has overall better performance and the reverse is true for $\alpha$ = 0.5 and 1. From these observations, we can conclude that for practical purposes, Vasicek-type estimator would be quite suitable for both normal and exponential distributions and for both choices of $m$.
4.1 Robustness of the estimators in the presence of outliers
@@
5 Estimation of Tsallis entropy based on progressive type-II censored data using $m$-spacings
@@
5.1 Empirical evaluation
We perform Monte Carlo simulations to assess the performance of the proposed estimator by examining its mean and variance. We consider different progressive censoring schemes for $n$ = 20 and $r$ = 10 and 15. We choose window size $m=[\sqrt{r}+0.5]$ and $\alpha$ = 1.10, 1.5 and 2. We mostly consider one-step progressive censoring schemes where all the surviving units are removed at one time during the experiment. We simulate 5000 progressively type-II censored samples from each of Exp(1) and N(0,1) distributions and calculate the average estimate (AE) and the variance of $\hat{T_{\alpha}}$. The obtained results are presented in Tables 3 and 4, respectively. In these tables, the notation aba*b refers to ‘$a$ is repeated $b$ times’. For instance, (05,10,04)(0*5,10,0*4) refers to $(0,0,0,0,0,10,0,0,0,0)$.
Table 3: Average Estimate and Variance of $\hat{T_{\alpha}}$ for Exp(1) distribution.
@@ 5.1 Empirical evaluation
Table 4: Average Estimate and Variance of $\hat{T_{\alpha}}$ for N(0,1) distribution.
@@
6 Estimation of Tsallis entropy through quantile approach
@@ 3. $K(\cdot)$ is symmetric about zero;
Proceeding along the same lines as in Theorems 6.1 and 6.2, we can establish the consistency and asymptotic normality of $\tilde{T}_{\alpha}(X)$ as well.
7 Empirical study
@@
and is denoted by Govindarajulu ($\mu,\sigma,\gamma$). We generate 5000 random samples from N(0,1) and Exp(1) distributions with sample size $n$ = 20, 50 and 100. We consider $\alpha$ = 0.5, 1.5, 2, 2.5 and 3 and then calculate the bias and the MSE of $T^{q}_{\alpha}(X)$. We use the Gaussian kernel and select bandwidth according to the normal reference rule $h_{n}=1.06\sigma n^{-\frac{1}{5}}$ (cf. Silverman, 2018). The obtained results are provided in Table 5. From this table, we observe that as the sample size increases, the bias and the MSE of $T^{q}_{\alpha}(X)$ for both distributions decrease. The estimator performs better when $\alpha$ increases. Also, $T^{q}_{\alpha}(X)$ performs better for N(0,1) than Exp(1) distribution. This may be due to the fact that we have chosen Gaussian Kernel and that the choice of $h_{n}$ is appropriate when data come from normal distribution.
Table 5: Bias and MSEs of $T^{q}_{\alpha}(X)$ for N(0,1) and Exp(1) distributions.
@@
There are some distributions that do not have tractable cdf, but have simple and closed form of qf, such as power-Pareto, skew lambda and Govindarajulu distributions. In these situations, it is advantageous to use $T^{q}_{\alpha}(X)$ to estimate Tsallis entropy. We simulate 5000 random samples from Govindarajulu distribution with parameters (0,0.75,0.25) and (0,0.25,0.75), for $n$ = 50, 100 and 200, and then calculate the bias and the MSE of $T^{q}_{\alpha}(X)$. We chose $\alpha$ = 0.15 and 0.75. Note that, for these models, Tsallis entropy does not exist for $\alpha>1$. We provide the obtained results in Table 6. The true value of Tsallis entropy are also included in the table. From this table, we observe that the estimator performs well for these models. Both Bias and MSE decrease as sample size increases and the performance of the estimator seems to be better when $\alpha$ is small.
Table 6: Bias and MSEs of $T^{q}_{\alpha}(X)$ for Govindarajulu distribution.
@@
8 Applications to Goodness-of-fit testing
@@
Table 7: Power of the tests for N(0,1) distribution.
@@ 8.1 Test for Normality
Table 8: Power of the tests for N(0,1) distribution.
@@ 8.2 Test for Exponentiality
The null hypothesis is rejected when the calculated value of $T>C_{2}$, where the critical point $C_{2}$ is obtained from the null distribution of $T$. For reference distribution, we consider Weibull (WE), gamma (GA) and log-normal (LN) distributions having monotone decreasing, increasing and non-monotone hazard functions. For monotone decreasing hazard alternatives, we take WE(0.5), GA(0.4) and LN(2). For monotone increasing hazard alternatives, we consider WE(2), GA(2) and GA(3), while for non-monotone hazard alternatives, we consider LN(0.6) and LN(1.2). The power of the tests so obtained are provided in Tables 9 and 10. From these tables, we find that proposed test performs better than the other two tests for monotone decreasing hazard alternative distributions.
Table 9: Power of the tests for Exp(1) distribution.
@@
Table 10: Power of the tests for Exp(1) distribution.
20(4) WE(0.5) 0.747 0.548 0.542 25(5) WE(0.5) 0.850 0.631 0.664
WE(2) 0.764 0.937 0.917 WE(2) 0.900 0.972 0.968
GA(0.4) 0.607 0.339 0.250 GA(0.4) 0.695 0.391 0.330
GA(2) 0.268 0.524 0.389 GA(2) 0.355 0.588 0.446
GA(3) 0.665 0.893 0.736 GA(3) 0.812 0.949 0.831
LN(0.6) 0.700 0.910 0.576 LN(0.6) 0.852 0.954 0.646
LN(1.2) 0.162 0.122 0.215 LN(1.2) 0.191 0.123 0.291
LN(2) 0.774 0.660 0.720 LN(2) 0.876 0.743 0.825
Exp(1) 0.051 0.054 0.047 Exp(1) 0.051 0.050 0.048
8.3 Goodness-of-fit Tests for Progressive Type-II Samples
@@
8.3.1 Exponentiality under Progressive Type-II Samples
For exponentiality testing, the test statistic is given by
where $\hat{\theta}=\frac{r}{\sum_{i=1}^{r}(1+R_{i})X_{i:r:n}}$ is the MLE for $\theta$ under PC-II censoring. We will reject exponentiality hypothesis if $\hat{T_{\alpha}}(F_{r,n},F)$ is large. We simulate 5000 random samples from Exp(1) distributions under various PC-II censoring schemes and obtain critical points at 10% significance level. Entropy based exponentiality tests under PC-II censoring perform better when the alternative distribution has increasing hazard function, see Balakrishnan et al. (2007). So, we consider WE(2) and GA(2) distributions for power calculation and compare the performance of the proposed test with the test proposed by Balakrishnan et al. (2007). Let us denote it by $T_{PC}$. We choose $\alpha$ = 2 and $m=[\sqrt{r}+0.5]$ for power computation and provide the results in Table 11. From the table we have observed that neither test outperforms the other for all the cases. Performance of both tests increases when sample size increases. Both tests perform better for WE(2) than for GA(2) alternative. The proposed test does attain the nominal level of significance. We mostly considered one step censoring schemes for illustrations. For other censoring schemes, the proposed tests outperforms $T_{PC}$ for both the alternative distributions.
Table 11: Power of exponentiality tests under PC-II samples at 10% significance level.
@@ 8.3.1 Exponentiality under Progressive Type-II Samples
8.4 Data Analyses